OpenStudy (anonymous):

Question Check The Question Was 25h + 40 = -15h - 8 25h + 15h = 40 +-80 40h = 120 h = 3 ?

OpenStudy (lalaly):

25h+40=-15h-80 25h+15h=-40-80 40h=-120 h=-3

OpenStudy (anonymous):

Check: 25*(-3)+40=-15*(-3)-80 -75+40=45-80 -35=-35